3 punkty drukarskie: 3 milimetry: 3 centymetry: 3 cale: 1 x wysokość litery x: 5.16667pt 5.16667pt ≈ 1 x rozmiar czcionki: 10.76656pt 11.74988pt

“rule[przesunięcie]–szerokość˝ –wysokość˝

Śliczny tekst pomiędzy krechami